OpenStudy (anonymous):

if a bus is going at the rate of 66km/h and the radius of wheel is 70cm then how many revolutions will be taken by bus in a minute.

OpenStudy (anonymous):

66 km/h = 18.3333 m/s wheel outside = 2*pi*r = 4.4 m 18.333/4.4*60 = 250 revs

OpenStudy (anonymous):

i got 25.3 something